# Thermonuclear Heating of Accreting Neutron Stars

### v1

@@ 1 Introduction
During our study of stationary accreting neutron star envelopes (Nava-Callejas et al., 2024a), we found that it may be possible to have accreting envelope models with a temperature inversion resulting in a negative luminosity $L_{\mathrm{b}}$ at its base and energy flowing into the crust from the envelope. Our purpose in the present work is to study such models in details and consider their implications for the heating of accreting neutron stars. As mentioned above, such solutions had already been found many years ago in fully time-dependent simulations (Hanawa & Fujimoto, 1984; Fujimoto et al., 1984, 1987) albeit with the simplification that they only included thermonuclear burning of H through the CNO cycle and no 4He burning. Later, Miralda-Escude et al. (1990) and Zdunik et al. (1992) presented similar but stationary models including the deep crustal heating, and also found solutions with temperature inversion within the envelope confirming the effect of enhanced core neutrino emission. More recently, Dohi et al. (2022) using their neutron star evolutionary code (Dohi et al., 2020) obtained comparable results, with the major improvements that they performed time-dependent calculations of the whole neutron star, envelope and interior, they included hydrogen burning through the rp-process (Wallace & Woosley, 1981), as well as the explosive helium burning, but they focused on the analysis of X-ray bursts. Their models, however, could only cover a few days of evolution of the star.
@@ 2 Stationary and Time-Dependent Accreting Envelopes
Figure 1: Examples of temperature profiles in stationary envelope models at three selected mass accretion rates, as marked in the upper part of the three panels. The (red) dashed lines denote the temperature $T_{\text{b}}$ at which $L_{\text{b}}=0$. Dots mark the point where nuclear energy generation reaches its maximum: at lower densities 1H burning dominates, at higher densities 4He burning makes a significant contribution for $\dot{M}=10^{-10}\,M_{\odot}$ yr-1 while at the higher $\dot{M}$ energy generation by the $\beta^{+}$ decays becomes dominant due to the large amount of heavy nuclei produced by the rp-process. Ample details of these phases of nuclear burning can be found in Nava-Callejas et al. (2024a).
@@
The calculations start at the photosphere with a chosen value of its effective temperature $T_{\mathrm{eff}}$, from which the surface density $\rho_{\mathrm{s}}$ and pressure $P_{\mathrm{s}}$ are deduced using an Eddington-type condition. More involved boundary conditions are possible, but they have little effect on the deeper layers as the initial condition is rapidly forgotten (Miralda-Escude et al., 1990; Kippenhahn et al., 2012). From this point on, the envelope equations are integrated inward until the density reaches a chosen boundary value $\rho_{\mathrm{b}}$. We chose $\rho_{\mathrm{b}}=10^{7}$ g cm-3 since the generation of nuclear energy past this value is negligible. The structure of the envelope in stationary condition depends on the mass accretion rate per unit area $\dot{m}=\dot{M}/4\pi R^{2}$ through the change in chemical composition when matter is being pushed from a depth $y$ to $y+dy$ and this change is proportional to the time $dt=dy/\dot{m}$ it takes for matter to move a depth $dy$. Moreover, when passing from the layer $y$ to $y+dy$ the luminosity $L(y)$ is reduced by the amount of nuclear and gravitational energy generated within this layer, starting from $L=4\pi R^{2}\sigma_{SB}T_{\mathrm{eff}}^{4}$ at the photosphere. When choosing a low enough $T_{\mathrm{eff}}$ it is possible for $L(y)$ to turn negative.
In Fig. 1 we illustrate the temperature profiles of three families of stationary envelopes, accreting a solar-like chemical composition, at three different rates. We consider a star of mass $M=1.4M_{\odot}$ and radius $R=11.56\,\mathrm{km}$ corresponding to an APR-core EOS star (Akmal et al., 1998). In all three cases we observe some envelopes where $T(y)$ reaches a maximum and then decreases, resulting in negative luminosities and a leakage of thermonuclear energy from the burning envelope into the neutron star interior. The dashed lines denote the temperature $T_{\text{b}}$ at which $L_{\text{b}}=0$, a critical value between envelopes with $L_{\text{b}}$ positive versus negative. Naturally this critical $T_{\text{b}}$ increases with increasing $\dot{M}$.
@@
[…]
@@
Figure 4: Comparison of the luminosity versus temperature, at four different accretion rates, of our stationary models, lines, and burst averages of our MESA time dependent models with boundary at $\rho_{b}=10^{7}$ g cm-3. Symbols have the same meaning as in Figure 3.
@@ 3 Stationary Accreting Envelopes as Boundary Conditions for Modeling Accreting Neutron Star Interiors
Figure 5: Left panel: envelope base luminosity, $L_{\mathrm{b}}$, versus mass accretion rate, $\dot{M}$, for several base temperatures, $T_{\mathrm{b}}$. Notice the change of $L_{\mathrm{b}}$ scale between the upper (positive $L_{\mathrm{b}}$) versus lower (negative $L_{\mathrm{b}}$) parts. Right panel: equivalent envelope base energy per baryon, $Q_{\mathrm{b}}$, versus mass accretion rate, $\dot{M}$, for several base temperatures, $T_{\mathrm{b}}$, corresponding to the $L_{\mathrm{b}}$ of the left panel with $L_{\mathrm{b}}<0$.
When modeling the thermal evolution of a neutron star one generally does not solve the thermal evolution equations down to the photosphere, but one rather introduces an outer boundary, located at a fixed density $\rho_{\mathrm{b}}$, or a fixed pressure $P_{\mathrm{b}}$, where an appropriate boundary condition is imposed. The latter is simply a relationship between the temperature, $T_{\mathrm{b}}$, and the luminosity, $L_{\mathrm{b}}$, at this outer boundary point (see, e.g., Appendix B in Page et al. 2004), i.e., a Cauchy boundary condition. Such relationship is obtained from envelope models as the ones presented for example by Gudmundsson et al. (1983) for envelopes made of pure iron or Potekhin et al. (1997), who generalized them to include accreted matter. Our stationary models presented in the previous section provide us with a further generalization to the case of accreting envelopes that automatically include in their structure the energy generated by the thermonuclear burning. For this purpose we employ our extensive net380 network (see Nava-Callejas et al. 2024a) and include general relativistic effects that were artificially excluded in the previous sections for comparison with MESA.
@@
As heating source we apply the deep crustal heating as described by Haensel & Zdunik (2008) that injects about 1.9 MeV baryon-1, distributed in the crust [^2] In the alternate model of Gusakov & Chugunov (2020) the energy released by the deep crustal heating is much smaller, of the order of 0.5 MeV (Potekhin et al., 2023). We considered this possibility and present the results in the “Supplementary Material” showing that it has only a very small effect on the impact of the crust heating from the envelope.. We also consider the effect of different amounts of shallow heating, $Q_{\mathrm{sh}}$, up to 3 MeV baryon-1 that are in the range of most values found necessary to interpret observations of crust relaxation in transient sources (see, e.g., Page & Reddy 2013; Degenaar et al. 2014; Turlione et al. 2015; Degenaar et al. 2015; Merritt et al. 2016; Parikh et al. 2018, 2019; Ootes et al. 2019; Degenaar et al. 2021; Page et al. 2022) and different densities, $\rho_{\mathrm{sh}}$, at which this energy is released, which we spread in the range $[\rho_{\mathrm{sh}},10\rho_{\mathrm{sh}}]$.
@@
Figure 7: Left panel: temperature profiles inside neutron star models persistently accreting at a rate of $10^{-9}\,M_{\odot}$ yr-1 and undergoing fast core neutrino cooling with $\log_{10}N_{8}^{\mathrm{fast}}\,[\mathrm{erg\,s}^{-1}]=38$. Dots at $\rho=10^{7}$ g cm-3 mark the interface between the stationary envelope and the interior: the plot shows interior profiles from NSCool connected to an envelope profile similar to the ones of Figure 1. Models A, B, and C, correspond to different amounts of shallow heating: 1 MeV baryon-1 deposited in the density range $[10^{8},10^{9}]$ g cm-3 (marked by a pastel red background) in A, and in the range $[10^{10},10^{11}]$ g cm-3 (marked by a pastel blue background) in B, while C has no shallow heating. Models plotted as continuous lines have a uniform $Q_{\mathrm{imp}}=10$ applied in the whole crust, while models plotted as dashed lines have no impurity added. The curves of the melting temperature, $T_{\mathrm{m}}$, corresponding to a Coulomb coupling parameter $\Gamma=180$, the neutron drip line, and the crust-core transition are also marked. Right panel: zoom-in of the frame of Figure 6 with the same values of $N^{\mathrm{fast}}=10^{38}$ erg s-1 and $Q_{\mathrm{sh}}=1$ MeV/baryon as in the left panel. The cyan color circles single out the six models depicted in the left panel.
@@
Notice that almost no models with a high accretion rate $\dot{M}>3\times 10^{-9}\,M_{\odot}$ yr-1, and only a small fraction of the ones with $\dot{M}=3\times 10^{-9}\,M_{\odot}$ yr-1, have a negative $Q_{\mathrm{b}}$, independently of the implemented $Q_{\mathrm{sh}}$: in all these situations we almost always have a strong heat flow from the crust into the envelope. It is interesting that these mass accretion rates are precisely the ones where observationally X-ray bursts are seen to be quenched.
@@
We here present a first step in applying our accreting envelopes to transiently accreting neutron stars. This is a direct extension of the method of the previous section with the intent to mimic the effect of recurrent outbursts of accretion separated by longer periods of quiescence. Considering the idealized case of a perfectly periodic system undergoing outbursts with constant mass accretion rate $\dot{M}_{\mathrm{ob}}$, this introduces at least two more parameters, which are the length $t_{\mathrm{ob}}$ of these outbursts and the recurrence time between one outburst and the next, $t_{\mathrm{rec}}$. For this first attempt, we will choose $t_{\mathrm{ob}}=$ 1 month and $t_{\mathrm{rec}}=10\,t_{\mathrm{ob}}$, i.e., a 10% duty cycle, which is roughly representative of several systems (Heinke et al., 2024)444 In the ”Suplementary Material” we present results for cases with longer accretion outburst and find very similar outcomes.. Due to the very high heat capacity of the neutron star core its temperature does not change appreciably during such short outbursts (Cumming et al., 2017) and the effect of many recurrent outbursts is to bring this temperature to a value such that, over a long time, the average energy losses by neutrinos and photons balance the average heating from the accretion (Colpi et al., 2001).
@@
As a complementary investigation we present in the “Supplementary Material” three cases where the outburst last 2 months, one year or one century but with very similar results.
@@
We summarize the results of these simulations in Figure 8 using the same presentation as in Figure 6, but emphasize that we now present the values of the average $\langle Q_{\mathrm{b}}\rangle$ for a one month-long accretion outburst. Overall, this figure has a similar appearance to Figure 6, but one clearly sees that there are many more cases with heat flowing from the envelope into the crust, i.e., with $\langle Q_{\mathrm{b}}\rangle<0$. When considering longer outbursts, but with the same duty cycle of 10%, very similar outcomes are obtained (see “Supplementary Material”). Longer outburst with larger duty cycles will become increasingly closer to the persistent cases.
@@
An animated version of this figure is available in the on-line version of this paper.
@@ 6 Discussion and Conclusions
A self-consistent evaluation of the possibility of having thermonuclear energy flowing into the stellar interior from the burning envelope requires modeling the whole star temperature evolution, i.e., the coupling between all possible energy sources and energy sinks (mostly neutrinos), over possibly long time scales, to determine in which cases the burning envelope is hotter or colder that the underling crust. Such models have already been described (Hanawa & Fujimoto, 1984; Fujimoto et al., 1984, 1987; Miralda-Escude et al., 1990; Zdunik et al., 1992; Dohi et al., 2022) under different conditions and assumptions. They clearly showed the possibility of an inward energy flux in the presence of very strong neutrino emission in the stellar core. Hanawa & Fujimoto (1984) and Dohi et al. (2022) followed in details the unstable burning, resulting in X-ray bursts, but could only study the evolution of the star for a few hours/days, as such calculations are extremely time consuming, while the other authors implemented stationary envelope models and could thus study the long term evolution of the star. To provide a stronger theoretical basis for the long term study of accreting neutron stars we showed, in Section 2, that stationary envelope models provide a reasonable first approximation to time-dependent envelope models when focusing on the amount of heat $Q_{\mathrm{b}}$ transferred between the crust and the envelope. When performing such evolutionary calculations the envelope simply provides a relationship between the temperature at the bottom of the envelope, $T_{\mathrm{b}}$, and the luminosity at that point, $L_{\mathrm{b}}$, or equivalently $Q_{\mathrm{b}}$ from Equation (1), that is used as an outer boundary condition for the modeling of the interior. We could provide an extensive set of such models, whose results are displayed in Figure 5, covering a wide range of mass accretion rates and boundary temperatures. The right panel of this figure shows that values of $Q_{\mathrm{b}}$ of more than one MeV baryon-1 are possible in the case of a cold crust.

### v3

@@ 1 Introduction
During our study of stationary accreting neutron star envelopes (Nava-Callejas et al., 2024a), we found that it may be possible to have accreting envelope models with a temperature inversion resulting in a negative luminosity $L_{\mathrm{b}}$ at its base and energy flowing into the crust from the envelope. Our purpose in the present work is to study such models in details and consider their implications for the heating of accreting neutron stars. As mentioned above, such solutions had already been found many years ago in fully time-dependent simulations (Hanawa & Fujimoto, 1984; Fujimoto et al., 1984, 1987) albeit with the simplification that they only included thermonuclear burning of H through the CNO cycle and no 4He burning. Later, Miralda-Escude et al. (1990) and Zdunik et al. (1992) presented similar but stationary models including the deep crustal heating, and also found solutions with temperature inversion within the envelope confirming the effect of enhanced core neutrino emission. More recently, Dohi et al. (2021, 2022) using their neutron star evolutionary code (Dohi et al., 2020) obtained comparable results, with the major improvements that they performed time-dependent calculations of the whole neutron star, envelope and interior, they included hydrogen burning through the rp-process (Wallace & Woosley, 1981), as well as the explosive helium burning, but they focused on the analysis of X-ray bursts. Their models, however, could only cover a few days of evolution of the star.
@@ 2 Stationary and Time-Dependent Accreting Envelopes
Figure 1: Examples of temperature profiles in stationary envelope models at three selected mass accretion rates, as marked in the upper part of the three panels. The horizontal lines denote the temperature $T_{\text{b}}$ at which $L_{\text{b}}=0$. Big dots mark the point where nuclear energy generation reaches its maximum: at lower densities 1H burning dominates, at higher densities 4He burning makes a significant contribution for $\dot{M}=10^{-10}\,M_{\odot}$ yr-1 while at the higher $\dot{M}$ energy generation by the $\beta^{+}$ decays becomes dominant due to the large amount of heavy nuclei produced by the rp-process. Ample details of these phases of nuclear burning can be found in Nava-Callejas et al. (2024a).
@@
The calculations start at the photosphere, at an optical depth of $2/3$, with a chosen value of its effective temperature $T_{\mathrm{eff}}$, from which the surface density $\rho_{\mathrm{s}}$ and pressure $P_{\mathrm{s}}$ are deduced using an Eddington-type condition. More involved boundary conditions are possible, but they have little effect on the deeper layers as the initial condition is rapidly forgotten (Miralda-Escude et al., 1990; Kippenhahn et al., 2012). From this point on, the envelope equations are integrated inward until the density reaches a chosen boundary value $\rho_{\mathrm{b}}$. We chose $\rho_{\mathrm{b}}=10^{7}$ g cm-3 since the generation of nuclear energy past this value is negligible. Such boundary condition, $\rho=10^{7}$ g cm-3, is a typical value in the literature of accreting neutron stars, both in stationary (e.g. Schatz et al. (1999)) or time-dependent simulations Heger et al. (2007). The structure of the envelope in stationary condition depends on the mass accretion rate per unit area $\dot{m}=\dot{M}/4\pi R^{2}$ through the change in chemical composition when matter is being pushed from a depth $y$ to $y+dy$ and this change is proportional to the time $dt=dy/\dot{m}$ it takes for matter to move a depth $dy$. Moreover, when passing from the layer $y$ to $y+dy$ the luminosity $L(y)$ is reduced by the amount of nuclear and gravitational energy generated within this layer, starting from $L=4\pi R^{2}\sigma_{SB}T_{\mathrm{eff}}^{4}$ at the photosphere. When choosing a low enough $T_{\mathrm{eff}}$ it is possible for $L(y)$ to turn negative.
In Fig. 1 we illustrate the temperature profiles of three families of stationary envelopes, accreting a solar-like chemical composition, at three different rates. We consider a star of mass $M=1.4M_{\odot}$ and radius $R=11.56\,\mathrm{km}$ - corresponding to an APR-core EOS star (Akmal et al., 1998). In all three cases we observe some envelopes where $T(y)$ reaches a maximum and then decreases, resulting in negative luminosities and a leakage of thermonuclear energy from the burning envelope into the neutron star interior. The horizontal lines denote the temperature $T_{\text{b}}$ at which $L_{\text{b}}=0$, a critical value between envelopes with $L_{\text{b}}$ positive versus negative. Naturally this critical $T_{\text{b}}$ increases with increasing $\dot{M}$.
@@
We display the temperature and luminosity averages in Figure 3 for the same four mass accretion rates, together with complementary models at different base luminosities and temperatures. For these averages we require a time interval: we employed the time interval between two successive peaks of effective temperature, i.e. a bursting cycle. The number of averages thus corresponds one-to-one with the number of simulated bursts and together they show a time evolutions similar to the stationary models curves. Unfortunately, reaching the stationary states would require much more computing time and would, anyway, only explore a very small part of the $L_{b}$$T_{b}$ parameter space. Nevertheless these results show that our stationary models appear to be a reasonable approximation.
@@
Figure 4: Comparison of the luminosity versus temperature, at four different accretion rates, of our stationary models, lines, and burst averages of our MESA time dependent models with boundary at $\rho_{b}=10^{7}$ g cm-3. Shaded regions are included to avoid ambiguities in the time dependent & independent comparison. Symbols have the same meaning as in Figure 3.
@@ 3 Stationary Accreting Envelopes as Boundary Conditions for Modeling Accreting Neutron Star Interiors
Figure 5: Left panel: envelope base luminosity, $L_{\mathrm{b}}$, versus mass accretion rate, $\dot{M}$, for several base temperatures, $T_{\mathrm{b}}$. Notice the change of $L_{\mathrm{b}}$ scale between the upper (positive $L_{\mathrm{b}}$) versus lower (negative $L_{\mathrm{b}}$) parts. Right panel: equivalent envelope base energy per baryon, $Q_{\mathrm{b}}$, versus mass accretion rate, $\dot{M}$, for several base temperatures, $T_{\mathrm{b}}$, corresponding to the $L_{\mathrm{b}}$ of the left panel with $L_{\mathrm{b}}<0$. A table of values of $L_{\mathrm{b}}$ interpolated on uniform grids of $T_{\mathrm{b}}$ and $\log_{10}\dot{M}$ from the ones used for this figure is available as a Table online.
When modeling the thermal evolution of a neutron star one generally does not solve the thermal evolution equations out to the photosphere, but one rather introduces an outer boundary, located at a fixed density $\rho_{\mathrm{b}}$, or a fixed pressure $P_{\mathrm{b}}$, where an appropriate boundary condition is imposed. The latter is simply a relationship between the temperature, $T_{\mathrm{b}}$, and the luminosity, $L_{\mathrm{b}}$, at this outer boundary point (see, e.g., Appendix B in Page et al. 2004), i.e., a Cauchy boundary condition. Such relationship is obtained from envelope models as the ones presented for example by Gudmundsson et al. (1983) for envelopes made of pure iron or Potekhin et al. (1997), who generalized them to include accreted matter. Our stationary models presented in the previous section provide us with a further generalization to the case of accreting envelopes that automatically include in their structure the energy generated by the thermonuclear burning. For this purpose we employ our extensive net380 network (see Nava-Callejas et al. 2024a) and include general relativistic effects that were artificially excluded in the previous sections for comparison with MESA.
@@
As heating source we apply the deep crustal heating as described by Haensel & Zdunik (2008) that injects about 1.9 MeV baryon-1, distributed in the crust [^2] In the alternate model of Gusakov & Chugunov (2020) the energy released by the deep crustal heating is much smaller, of the order of 0.5 MeV (Potekhin et al., 2023). We considered this possibility but noticed it has only a very small effect on the impact of the crust heating from the envelope.. We also consider the effect of different amounts of shallow heating, $Q_{\mathrm{sh}}$, up to 3 MeV baryon-1 that are in the range of most values found necessary to interpret observations of crust relaxation in transient sources (see, e.g., Page & Reddy 2013; Degenaar et al. 2014; Turlione et al. 2015; Degenaar et al. 2015; Merritt et al. 2016; Parikh et al. 2018, 2019; Ootes et al. 2019; Degenaar et al. 2021; Page et al. 2022) and different densities, $\rho_{\mathrm{sh}}$, at which this energy is released, which we spread in the range $\rho_{\mathrm{sh}},10\rho_{\mathrm{sh}}]$.
@@
Figure 7: Left panel: temperature profiles inside neutron star models persistently accreting at a rate of $10^{-9}\,M_{\odot}$ yr-1 and undergoing fast core neutrino cooling with $\log_{10}N_{8}^{\mathrm{fast}}\,[\mathrm{erg\,s}^{-1}]=38$. Dots at $\rho=10^{7}$ g cm-3 mark the interface between the stationary envelope and the interior: the plot shows interior profiles from NSCool connected to an envelope profile similar to the ones of Figure 1. Models A, B, and C, correspond to different amounts of shallow heating: 1 MeV baryon-1 deposited in the density range $10^{8},10^{9}]$ g cm-3 (marked by a pastel red background) in A, and in the range $10^{10},10^{11}]$ g cm-3 (marked by a pastel blue background) in B, while C has no shallow heating. Models plotted as continuous lines have a uniform $Q_{\mathrm{imp}}=10$ applied in the whole crust, while models plotted as dashed lines have no impurity added. The curves of the melting temperature, $T_{\mathrm{m}}$, corresponding to a Coulomb coupling parameter $\Gamma=180$, the neutron drip line, and the crust-core transition are also marked. Right panel: zoom-in of the frame of Figure 6 with the same values of $N^{\mathrm{fast}}=10^{38}$ erg s-1 and $Q_{\mathrm{sh}}=1$ MeV/baryon as in the left panel. The cyan color circles single out the six models depicted in the left panel.
@@
Notice that almost no models with a high accretion rate $\dot{M}>3\times 10^{-9}\,M_{\odot}$ yr-1, and only a small fraction of the ones with $\dot{M}=3\times 10^{-9}\,M_{\odot}$ yr-1, have a negative $Q_{\mathrm{b}}$, independently of the implemented $Q_{\mathrm{sh}}$: in all these situations we almost always have a strong heat flow from the crust into the envelope. It is interesting that these mass accretion rates are precisely the ones where observationally X-ray bursts are seen to be quenched (e.g. Cornelisse et al. (2003); Galloway et al. (2008); Galloway & Keek (2021)).
@@
We here present a first step in applying our accreting envelopes to transiently accreting neutron stars. This is a direct extension of the method of the previous section with the intent to mimic the effect of recurrent outbursts of accretion separated by longer periods of quiescence. Considering the idealized case of a perfectly periodic system undergoing outbursts with constant mass accretion rate $\dot{M}_{\mathrm{ob}}$, this introduces at least two more parameters, which are the length $t_{\mathrm{ob}}$ of these outbursts and the recurrence time between one outburst and the next, $t_{\mathrm{rec}}$. For this first attempt, we will choose $t_{\mathrm{ob}}=$ 1 month and $t_{\mathrm{rec}}=10\,t_{\mathrm{ob}}$, i.e., a 10% duty cycle, which is roughly representative of several systems (Heinke et al., 2024). Due to the very high heat capacity of the neutron star core its temperature does not change appreciably during such short outbursts (Cumming et al., 2017) and the effect of many recurrent outbursts is to bring this temperature to a value such that, over a long time, the average energy losses by neutrinos and photons balance the average heating from the accretion (Colpi et al., 2001).
@@
We summarize the results of these simulations in Figure 8 using the same presentation as in Figure 6, but emphasize that we now present the values of the average $\langle Q_{\mathrm{b}}\rangle$ for a one month-long accretion outburst. Overall, this figure has a similar appearance to Figure 6, but one clearly sees that there are many more cases with heat flowing from the envelope into the crust, i.e., with $\langle Q_{\mathrm{b}}\rangle<0$. When considering longer outbursts, but with the same duty cycle of 10%, very similar outcomes are obtained. Longer outburst with larger duty cycles will become increasingly closer to the persistent cases.
@@
An animated version of this figure is available in the on-line version of this paper: it presents the time evolution, from day -1 to day +300, of the temperature profiles displayed in the four panels of the figure.
@@ 6 Discussion and Conclusions
A self-consistent evaluation of the possibility of having thermonuclear energy flowing into the stellar interior from the burning envelope requires modeling the whole star temperature evolution, i.e., the coupling between all possible energy sources and energy sinks (mostly neutrinos), over possibly long time scales, to determine in which cases the burning envelope is hotter or colder that the underling crust. Such models have already been described (Hanawa & Fujimoto, 1984; Fujimoto et al., 1984, 1987; Miralda-Escude et al., 1990; Zdunik et al., 1992; Dohi et al., 2021, 2022) under different conditions and assumptions. They clearly showed the possibility of an inward energy flux in the presence of very strong neutrino emission in the stellar core. Hanawa & Fujimoto (1984) and Dohi et al. (2021, 2022) followed in details the unstable burning, resulting in X-ray bursts, but could only study the evolution of the star for a few hours/days, as such calculations are extremely time consuming, while the other authors implemented stationary envelope models and could thus study the long term evolution of the star. To provide a stronger theoretical basis for the long term study of accreting neutron stars we showed, in Section 2, that stationary envelope models provide a reasonable first approximation to time-dependent envelope models when focusing on the amount of heat $Q_{\mathrm{b}}$ transferred between the crust and the envelope. When performing such evolutionary calculations the envelope simply provides a relationship between the temperature at the bottom of the envelope, $T_{\mathrm{b}}$, and the luminosity at that point, $L_{\mathrm{b}}$, or equivalently $Q_{\mathrm{b}}$ from Equation (1), that is used as an outer boundary condition for the modeling of the interior. We could provide an extensive set of such models, whose results are displayed in Figure 5, covering a wide range of mass accretion rates and boundary temperatures. The right panel of this figure shows that values of $Q_{\mathrm{b}}$ of more than one MeV baryon-1 are possible in the case of a cold crust.